# Meta-RL Induces Exploration in Language Agents

### v1

@@ 1 Introduction
Recent advances in large language models (LLMs) have shifted from building conversational systems to decision-making agents capable of reasoning about and interacting with their environments (yao2023react; shinn2023reflexion; wang2025ragen; feng2025group). To accomplish the goal, language agents operate in multi-turn, textual observation-action loops, and must adapt quickly using the memory across turns. Central to such adaptation is exploration, which allows agents to test uncertain actions, acquire new knowledge, and avoid premature convergence on suboptimal strategies. However, unlike humans that can explore systematically and make fast adaptation in new environments (wilson2014humans), LLM agents do not robustly engage in exploration without substantial interventions (krishnamurthy2024can).
Recent works have begun to address this limitation by guiding LLMs toward exploratory behaviors at test time. For example, tajwar2025training train models offline to distill exploration strategies from trajectories across diverse environments, while gandhi2024stream induce such strategies from offline search traces. setlur2025e3 train models to learn to explore in-context as a better way of spending test-time compute (snell2025scaling). However, these works either focus on single-turn non-agentic reasoning problems, or rely on offline data which limits them to imitation rather than active exploration.
In this work, we take a step toward agents that can actively explore their environment, gather feedback, and leverage this experience for more effective exploitation. Since multi-turn tasks often have a sparse success signal after an episode, we consider a multi-episode regime (shinn2023reflexion) where an episode is the unit of exploration and exploitation. Balancing exploration and exploitation can then be naturally formulated as a cross-episode reinforcement learning (RL) framework. Training across many similar but different environments under this framework leads to the meta reinforcement learning (Meta-RL) (duan2016rl; wang2016learning; bauer2023human; Beck_2025), where the agent is forced to discover general strategies that work in unseen and potentially harder environments.
@@
We evaluate LaMer on four challenging long-horizon tasks: Sokoban (racaniere2017imagination), MineSweeper (Li.2023.Minesweeper), Webshop (yao2022webshop) and ALFWorld (shridhar2020alfworld). Using Qwen3-4B (yang2025qwen3), we demonstrate that LaMer consistently outperforms prompting and RL baselines across all environments. In addition, we observe that the trained model learns to balance between exploration and exploitation, resulting in improved test-time scaling performance. In particular, LaMer adapts the trained policy at test time, with $11/14/19\%$ absolute performance gains on Sokoban, MineSweeper and Webshop, respectively, over the RL. Furthermore, we show that the LaMer trained model generalizes better to harder and out-of-distribution tasks. In summary, LaMer takes a step toward autonomous agents that can actively act to uncover information and improve their decision-making in the new environments.
@@ 2 Related Work
LLM-as-agent. As LLMs become increasingly capable of reasoning about complex scenarios (wei2022chain), there is a growing interest in making them decision-making autonomous agents. Earlier works rely on prompting frozen LLMs (yao2023react; shinn2023reflexion; park2023generative; wang2024voyager; autogpt). ReAct (yao2023react) prompts LLMs with in-context examples to generate both textual actions and reasoning thoughts. Later, Reflexion (shinn2023reflexion) extends this principle to the multi-episode setting, where the agent verbally reflects on the last episode and maintains their own reflection buffer for the next episodes. More recent research trains LLM agents through designing advanced RL algorithms (wang2025ragen; feng2025group) for multi-turn interactions, or supervised fine-tuning on generated interaction trajectories across diverse tasks (tajwar2025training). The evaluation of LLM agents also poses challenges because of fully verbal interactions with the environments. Recent benchmarks span a wide range of domains, including text-based embodied environments (shridhar2020alfworld), e-commerce website (yao2022webshop), bandits (nie2024evolve), classic games (park2025orak; Li.2023.Minesweeper) and other tasks (liu2024agentbench; nathani2025mlgymnewframeworkbenchmark). For a more comprehensive overview of these efforts, we refer readers to recent surveys (wang2024survey; zhang2025landscape).
@@
Meta reinforcement learning (Meta-RL) (Beck_2025) focuses on “learning to reinforcement learn” in order to rapidly adapt to new environments. Similar to meta-learning (thrun1998learning; hospedales2021meta), it involves an inner-loop that represents an RL algorithm (i.e., an adaptation strategy) by itself, together with an outer-loop that updates the meta-parameters so that the inner loop becomes more effective across many tasks. By training across many tasks, the outer loop forces the agent to learn the exploration strategy necessary to solve the tasks, while the inner loop enables the agent adapt quickly based on the exploration. Depending on how the inner loop is done, there are in-context methods and gradient-based methods. For example, duan2016rl; wang2016learning; stadie2018some represent the inner-loop as a history-dependent policy parametrized by an RNN, the adaptation is thus done ‘in-context’ through gathering more information stored in the memory states. On the other hand, finn2017model leverages a gradient-based approach in which the inner adapts a general meta-policy learned by the outer-loop. Our work lies in the former category, where the adaptation occurs entirely in-context at test time, naturally leveraging LLMs’ in-context learning abilities.
@@
Test-time compute. A Meta-RL framework in LaMer can be considered as amortizing the test-time compute by training tasks in a multi-episode manner rather than a single episode. In this way, the learned in-context policy adaptation balances exploration and exploitation for a fast adaptation at test time. This is essentially a better way of spending test-time compute (snell2025scaling; muennighoff2025s1simpletesttimescaling; wu2025inferencescalinglawsempirical; setlur2025e3). In our experiments, we match the training compute budget between an RL and a Meta-RL baseline, and show that meta-RL encourages superior test-time scaling behavior (through pass@k). qu2025optimizing similarly relates meta-RL to test-time compute, but they are limited to single-turn problems of mathematical reasoning, without leveraging the interactive feedback from the environment.
@@
Reasoning in LLMs. More broadly, this work relates to reasoning in LLMs, because language agents must use reasoning as part of their decision-making. A large bulk of recent work on LLM reasoning has focused on more advanced prompting (wei2022chain; yao2023tree), post-training (cobbe2021training; luong2024reftreasoningreinforcedfinetuning; shao2024deepseekmath; deepseekai2025deepseekr1incentivizingreasoningcapability) or bootstrapping (zelikman2022star) against verifiers or reward models, inducing structured search behavior (gandhi2024stream; moon2024guidedstreamsearchlearning), or reflecting on previous answers (kumar2024training; xiong2025selfrewardingcorrectionmathematicalreasoning; qu2024recursive), etc. Most of these works focus on single-turn math (hendrycks2021measuring; cobbe2021training) and coding (chen2021codex; hendrycks2021measuringcodingchallengecompetence) problems, while we target multi-turn agentic environments where environment feedback is available after every action and at the end of the episode.
@@ 3 Preliminaries
where $\gamma_{\text{step}}\in[0,1]$ is the discount factor. Recent works (wang2025ragen; feng2025group) have shown that RL training has enabled LLM agents to interact with the environment and solve multi-turn tasks. However, such agents often learn a fixed policy during training and struggle to actively explore and adapt their behavior to the tasks at test time (nie2024evolve).
Meta-RL. Conversely, by training on a distribution of tasks, Meta-RL  (duan2016rl; wang2016learning; bauer2023human; Beck_2025) encourages exploration because it optimizes meta-parameters, such that the agent can solve new tasks quickly. In our case, the meta-parameters are the parameters of the LLM. This necessitates the agent to learn general exploration-exploitation strategies suitable for the task distribution trained on. For example, for most navigation tasks in partially observable environments, the optimal strategy is to gather the environment information and locate the target during the first episode, then reach the target as efficiently as possible in the second episode. This explore-then-exploit strategy implemented by the agent is itself an RL algorithm, where the policy learned at the meta-level encodes how to adaptively switch between information-gathering and reward-maximizing behaviors depending on the stage of interaction with a new task. For LLM agents operating in multi-turn tasks, the policy can be in context (i.e., without parameter updates at test time), naturally leveraging the in-context capability of LLMs.
@@ 4 LaMer: A Meta-RL Framework for LLM agents
In the Meta-RL, policy adaptation is the inner loop of the learning process of an LLM agent. Therefore, a flexible and efficient adaptation mechanism plays an important role during training and methods like gradient descent (finn2017model) might be too expensive, especially for LLMs. In LaMer, we propose a self-reflection based strategy(shinn2023reflexion) to adapt the policy in-context (brown2020language; laskin2023incontext). Specifically, after each episode finishes, we prompt the agent to generate the textual reflection on the previous attempt, providing specific feedback and plan to guide the next episode (see Appendix A for the used prompt). The policy is therefore updated through modifying the context, $\pi^{(n)}_{\theta}(\cdot)=\pi_{\theta}(\cdot|\mathcal{H}^{(n)})$ where $\mathcal{H}^{(n)}$ denotes the inter-episode memory that contains both the history trajectories and reflections. Importantly, the self-reflection step is also explicitly trained in LaMer using the reward obtained in the next episode. Note that the content $\mathcal{H}^{(n)}$ can be adjusted according to the predefined memory buffer to reduce the context length and improve the efficiency. By default, we retain both history and reflection in $\mathcal{H}^{(n)}$, and provide an ablation study in Section 6.2.
@@
where $A_{t}^{(n)}$ is the advantage estimation derived from $G_{t}^{(n)}$. The framework is compatible with widely used optimizers such PPO (schulman2017proximal) and critic-free approaches such as GRPO (shao2024deepseekmath) and GiGPO (feng2025group).
@@ 5.1 Experimental Setup
We evaluate LaMer on four challenging and diverse environments: Sokoban (racaniere2017imagination), MineSweeper (Li.2023.Minesweeper), Webshop (yao2022webshop) and ALFWorld (shridhar2020alfworld). Among them, Sokoban is a classic grid-based game on planning where the environment is fully observable. In comparison, the environments of MineSweeper, ALFWorld and Webshop are partially observable, requiring the agent to explore and plan under uncertainty to finish the task. Specifically, MineSweeper is a board game about logical deduction on hidden cells. Webshop simulates realistic web-based shopping tasks, and ALFWorld provides text-based embodied environments. We provide the detailed explanation and prompts of the environment in Appendix A. All the experiments are conducted with the text modality, though the proposed method can be naturally applied to multimodal environments.
@@
We use Qwen3-4B (yang2025qwen3) as our base model for all the experiments. To improve rollout efficiency in agentic loops, we use the non-thinking mode during trajectory generation. Additionally, we validate our method on Llama3.1-8B-Instruct (grattafiori2024llama3herdmodels), see Appendix D.1 for the results. For the Meta-RL setting, we use $\gamma_{\text{traj}}=0.6$ as the default trajectory discount factor and explore its influence in the ablation study. We use GiGPO as the default optimization algorithm for all the environments with LaMer. Importantly, for Meta-RL training we sample $N=3$ episodes and set group size to $8$ for each task. To ensure a fair comparison, we use a group size of $24$ in standard RL training, yielding the same number of trajectories used for each gradient update step. All other hyperparameters and configurations are kept identical across RL and Meta-RL for a fair comparison and are provided in Appendix C. Our code is publicly available at https://github.com/mlbio-epfl/LaMer.
@@ 5.2 Performance comparison
We compare the performance of the LaMer framework with prompting-based baselines (Zero-shot, ReAct (yao2023react; shinn2023reflexion)), and RL methods (PPO (schulman2017proximal), RLOO (ahmadian2024back), GRPO (shao2024deepseekmath), and GiGPO (feng2025group)) across three environments: Sokoban, MineSweeper, and Webshop. For each method, we report the success rates under 1, 2, and 3 attempts (i.e., pass@1, pass@2, and pass@3, respectively). The results are summarized in Table 1.
@@ 5.4 Generalization to unseen tasks
We further study the ability LaMer and alternative methods to generalize out-of-distribution. For this experiment, we use the ALFWorld environment (shridhar2020alfworld). As a text-based embodied environment, ALFWorld contains 6 categories of common household activities: Pick and Place (Pick), Examine in Light (Look), Clean and Place (Clean), Heat and Place (Heat), Cool and Place (Cool), and Pick Two and Place (Pick2). We use the tasks of Pick, Look, Clean and Heat as in-distribution tasks and use Cool and Pick2 as out-of-distribution tasks. We train LaMer and alternative baselines with instances from in-distribution tasks and then evaluate the model on both in-distribution tasks (with held-out test set), and the examples of out-of-distribution tasks. The results are shown in Table 2. As we can see, RL trained model generally performs well on in-distribution tasks and outperforms prompting-based methods by achieving more than 20% improvement on Look, Clean and Heat. However, on out-of-distribution tasks, Cool and Pick2, it only obtains 58.1% and 36.0% success rate. LaMer with Meta-RL consistently outperforms RL on both in-distribution and out-of-distribution tasks, with a notable performance gap on out-of-distribution tasks. In particular, our LaMer framework achieves 23% performance gains on Cool and around 14% on Pick2. Overall, these results suggest that on ALFWorld, Meta-RL trained model could generalize better to out-of-distribution tasks compared to the RL trained model.

### v2

@@ 1 Introduction
Recent advances in large language models (LLMs) have shifted from building conversational systems to decision-making agents capable of reasoning about and interacting with their environments (Yao et al., 2023b; Shinn et al., 2023; Wang et al., 2025; Feng et al., 2025). To accomplish the goal, language agents operate in multi-turn, textual observation-action loops, and must adapt quickly using the memory across turns. Central to such adaptation is exploration, which allows agents to test uncertain actions, acquire new knowledge, and avoid premature convergence on suboptimal strategies. However, unlike humans that can explore systematically and make fast adaptation in new environments (Wilson et al., 2014), LLM agents do not robustly engage in exploration without substantial interventions (Krishnamurthy et al., 2024).
Recent works have begun to address this limitation by guiding LLMs toward exploratory behaviors at test time. For example, Tajwar et al. (2025) train models offline to distill exploration strategies from trajectories across diverse environments, while Gandhi et al. (2024) induce such strategies from offline search traces. Setlur et al. (2025) train models to learn to explore in-context as a better way of spending test-time compute (Snell et al., 2025). However, these works either focus on single-turn non-agentic reasoning problems, or rely on offline data which limits them to imitation rather than active exploration.
In this work, we take a step toward agents that can actively explore their environment, gather feedback, and leverage this experience for more effective exploitation. Since multi-turn tasks often have a sparse success signal after an episode, we consider a multi-episode regime (Shinn et al., 2023) where an episode is the unit of exploration and exploitation. Balancing exploration and exploitation can then be naturally formulated as a cross-episode reinforcement learning (RL) framework. Training across many similar but different environments under this framework leads to the meta reinforcement learning (Meta-RL) (Duan et al., 2016; Wang et al., 2016; Bauer et al., 2023; Beck et al., 2025), where the agent is forced to discover general strategies that work in unseen and potentially harder environments.
@@
We evaluate LaMer on four challenging long-horizon tasks: Sokoban (Racanière et al., 2017), MineSweeper (Li et al., 2024), Webshop (Yao et al., 2022) and ALFWorld (Shridhar et al., 2021). Using Qwen3-4B (Yang et al., 2025), we demonstrate that LaMer consistently outperforms prompting and RL baselines across all environments. In addition, we observe that the trained model learns to balance between exploration and exploitation, resulting in improved test-time scaling performance. In particular, LaMer adapts the trained policy at test time, with $11/14/19\%$ absolute performance gains on Sokoban, MineSweeper and Webshop, respectively, over the RL. Furthermore, we show that the LaMer trained model generalizes better to harder and out-of-distribution tasks. In summary, LaMer takes a step toward autonomous agents that can actively act to uncover information and improve their decision-making in the new environments.
@@ 2 Related Work
LLM-as-agent. As LLMs become increasingly capable of reasoning about complex scenarios (Wei et al., 2022), there is a growing interest in making them decision-making autonomous agents. Earlier works rely on prompting frozen LLMs (Yao et al., 2023b; Shinn et al., 2023; Park et al., 2023; Wang et al., 2024a; AutoGPT, 2023). ReAct (Yao et al., 2023b) prompts LLMs with in-context examples to generate both textual actions and reasoning thoughts. Later, Reflexion (Shinn et al., 2023) extends this principle to the multi-episode setting, where the agent verbally reflects on the last episode and maintains their own reflection buffer for the next episodes. More recent research trains LLM agents through designing advanced RL algorithms (Wang et al., 2025; Feng et al., 2025) for multi-turn interactions, or supervised fine-tuning on generated interaction trajectories across diverse tasks (Tajwar et al., 2025). The evaluation of LLM agents also poses challenges because of fully verbal interactions with the environments. Recent benchmarks span a wide range of domains, including text-based embodied environments (Shridhar et al., 2021), e-commerce website (Yao et al., 2022), bandits (Nie et al., 2024), classic games (Park et al., 2025; Li et al., 2024) and other tasks (Liu et al., 2024; Nathani et al., 2025). For a more comprehensive overview of these efforts, we refer readers to recent surveys (Wang et al., 2024b; Zhang et al., 2025).
@@
Meta reinforcement learning (Meta-RL) (Beck et al., 2025) focuses on “learning to reinforcement learn” in order to rapidly adapt to new environments. Similar to meta-learning (Thrun and Pratt, 1998; Hospedales et al., 2021), it involves an inner-loop that represents an RL algorithm (i.e., an adaptation strategy) by itself, together with an outer-loop that updates the meta-parameters so that the inner loop becomes more effective across many tasks. By training across many tasks, the outer loop forces the agent to learn the exploration strategy necessary to solve the tasks, while the inner loop enables the agent adapt quickly based on the exploration. Depending on how the inner loop is done, there are in-context methods and gradient-based methods. For example, Duan et al. (2016); Wang et al. (2016); Stadie et al. (2018) represent the inner-loop as a history-dependent policy parametrized by an RNN, the adaptation is thus done ‘in-context’ through gathering more information stored in the memory states. On the other hand, Finn et al. (2017) leverages a gradient-based approach in which the inner adapts a general meta-policy learned by the outer-loop. Our work lies in the former category, where the adaptation occurs entirely in-context at test time, naturally leveraging LLMs’ in-context learning abilities.
@@
Test-time compute. A Meta-RL framework in LaMer can be considered as amortizing the test-time compute by training tasks in a multi-episode manner rather than a single episode. In this way, the learned in-context policy adaptation balances exploration and exploitation for a fast adaptation at test time. This is essentially a better way of spending test-time compute (Snell et al., 2025; Muennighoff et al., 2025; Wu et al., 2025; Setlur et al., 2025). In our experiments, we match the training compute budget between an RL and a Meta-RL baseline, and show that meta-RL encourages superior test-time scaling behavior (through pass@k). Qu et al. (2025) similarly relates meta-RL to test-time compute, but they are limited to single-turn problems of mathematical reasoning, without leveraging the interactive feedback from the environment.
@@
Reasoning in LLMs. More broadly, this work relates to reasoning in LLMs, because language agents must use reasoning as part of their decision-making. A large bulk of recent work on LLM reasoning has focused on more advanced prompting (Wei et al., 2022; Yao et al., 2023a), post-training (Cobbe et al., 2021; Luong et al., 2024; Shao et al., 2024; DeepSeek-AI et al., 2025) or bootstrapping (Zelikman et al., 2022) against verifiers or reward models, inducing structured search behavior (Gandhi et al., 2024; Moon et al., 2024), or reflecting on previous answers (Kumar et al., 2024; Xiong et al., 2025; Qu et al., 2024), etc. Most of these works focus on single-turn math (Hendrycks et al., 2021b; Cobbe et al., 2021) and coding (Chen et al., 2021; Hendrycks et al., 2021a) problems, while we target multi-turn agentic environments where environment feedback is available after every action and at the end of the episode.
@@ 3 Preliminaries
where $\gamma_{\text{step}}\in[0,1]$ is the discount factor. Recent works (Wang et al., 2025; Feng et al., 2025) have shown that RL training has enabled LLM agents to interact with the environment and solve multi-turn tasks. However, such agents often learn a fixed policy during training and struggle to actively explore and adapt their behavior to the tasks at test time (Nie et al., 2024).
Meta-RL. Conversely, by training on a distribution of tasks, Meta-RL  (Duan et al., 2016; Wang et al., 2016; Bauer et al., 2023; Beck et al., 2025) encourages exploration because it optimizes meta-parameters, such that the agent can solve new tasks quickly. In our case, the meta-parameters are the parameters of the LLM. This necessitates the agent to learn general exploration-exploitation strategies suitable for the task distribution trained on. For example, for most navigation tasks in partially observable environments, the optimal strategy is to gather the environment information and locate the target during the first episode, then reach the target as efficiently as possible in the second episode. This explore-then-exploit strategy implemented by the agent is itself an RL algorithm, where the policy learned at the meta-level encodes how to adaptively switch between information-gathering and reward-maximizing behaviors depending on the stage of interaction with a new task. For LLM agents operating in multi-turn tasks, the policy can be in context (i.e., without parameter updates at test time), naturally leveraging the in-context capability of LLMs.
@@ 4 LaMer: A Meta-RL Framework for LLM agents
In the Meta-RL, policy adaptation is the inner loop of the learning process of an LLM agent. Therefore, a flexible and efficient adaptation mechanism plays an important role during training and methods like gradient descent (Finn et al., 2017) might be too expensive, especially for LLMs. In LaMer, we propose a self-reflection based strategy(Shinn et al., 2023) to adapt the policy in-context (Brown et al., 2020; Laskin et al., 2023). Specifically, after each episode finishes, we prompt the agent to generate the textual reflection on the previous attempt, providing specific feedback and plan to guide the next episode (see Appendix A for the used prompt). The policy is therefore updated through modifying the context, $\pi^{(n)}_{\theta}(\cdot)=\pi_{\theta}(\cdot|\mathcal{H}^{(n)})$ where $\mathcal{H}^{(n)}$ denotes the inter-episode memory that contains both the history trajectories and reflections. Importantly, the self-reflection step is also explicitly trained in LaMer using the reward obtained in the next episode. Note that the content $\mathcal{H}^{(n)}$ can be adjusted according to the predefined memory buffer to reduce the context length and improve the efficiency. By default, we retain both history and reflection in $\mathcal{H}^{(n)}$, and provide an ablation study in Section 6.2.
@@
where $A_{t}^{(n)}$ is the advantage estimation derived from $G_{t}^{(n)}$. The framework is compatible with widely used optimizers such PPO (Schulman et al., 2017) and critic-free approaches such as GRPO (Shao et al., 2024) and GiGPO (Feng et al., 2025).
@@ 5.1 Experimental Setup
We evaluate LaMer on four challenging and diverse environments: Sokoban (Racanière et al., 2017), MineSweeper (Li et al., 2024), Webshop (Yao et al., 2022) and ALFWorld (Shridhar et al., 2021). Among them, Sokoban is a classic grid-based game on planning where the environment is fully observable. In comparison, the environments of MineSweeper, ALFWorld and Webshop are partially observable, requiring the agent to explore and plan under uncertainty to finish the task. Specifically, MineSweeper is a board game about logical deduction on hidden cells. Webshop simulates realistic web-based shopping tasks, and ALFWorld provides text-based embodied environments. We provide the detailed explanation and prompts of the environment in Appendix A. All the experiments are conducted with the text modality, though the proposed method can be naturally applied to multimodal environments.
@@
We use Qwen3-4B (Yang et al., 2025) as our base model for all the experiments. To improve rollout efficiency in agentic loops, we use the non-thinking mode during trajectory generation. Additionally, we validate our method on Llama3.1-8B-Instruct (Grattafiori et al., 2024), see Appendix D.1 for the results. For the Meta-RL setting, we use $\gamma_{\text{traj}}=0.6$ as the default trajectory discount factor and explore its influence in the ablation study. We use GiGPO as the default optimization algorithm for all the environments with LaMer. Importantly, for Meta-RL training we sample $N=3$ episodes and set group size to $8$ for each task. To ensure a fair comparison, we use a group size of $24$ in standard RL training, yielding the same number of trajectories used for each gradient update step. All other hyperparameters and configurations are kept identical across RL and Meta-RL for a fair comparison and are provided in Appendix C. Our code is publicly available at https://github.com/mlbio-epfl/LaMer.
@@ 5.2 Performance comparison
We compare the performance of the LaMer framework with prompting-based baselines (Zero-shot, ReAct (Yao et al., 2023b; Shinn et al., 2023)), and RL methods (PPO (Schulman et al., 2017), RLOO (Ahmadian et al., 2024), GRPO (Shao et al., 2024), and GiGPO (Feng et al., 2025)) across three environments: Sokoban, MineSweeper, and Webshop. For each method, we report the success rates under 1, 2, and 3 attempts (i.e., pass@1, pass@2, and pass@3, respectively). The results are summarized in Table 1.
@@ 5.4 Generalization to unseen tasks
We further study the ability LaMer and alternative methods to generalize out-of-distribution. For this experiment, we use the ALFWorld environment (Shridhar et al., 2021). As a text-based embodied environment, ALFWorld contains 6 categories of common household activities: Pick and Place (Pick), Examine in Light (Look), Clean and Place (Clean), Heat and Place (Heat), Cool and Place (Cool), and Pick Two and Place (Pick2). We use the tasks of Pick, Look, Clean and Heat as in-distribution tasks and use Cool and Pick2 as out-of-distribution tasks. We train LaMer and alternative baselines with instances from in-distribution tasks and then evaluate the model on both in-distribution tasks (with held-out test set), and the examples of out-of-distribution tasks. The results are shown in Table 2. As we can see, RL trained model generally performs well on in-distribution tasks and outperforms prompting-based methods by achieving more than 20% improvement on Look, Clean and Heat. However, on out-of-distribution tasks, Cool and Pick2, it only obtains 58.1% and 36.0% success rate. LaMer with Meta-RL consistently outperforms RL on both in-distribution and out-of-distribution tasks, with a notable performance gap on out-of-distribution tasks. In particular, our LaMer framework achieves 23% performance gains on Cool and around 14% on Pick2. Overall, these results suggest that on ALFWorld, Meta-RL trained model could generalize better to out-of-distribution tasks compared to the RL trained model.